Suppose y varies inversely with x. Write an equation for the inverse variation.
y=6 when x=3

Answer by ortarico(23)

To show this:
y = k/x, where k is the constant
if x = 3 and y = 6
:. 6 = k/3
k = 18
--------------

y = 18/x
:. Equation for this is: xy = 18